OpenStudy (anonymous):

Identify the vertex for the graph of y = -2x2 - 12x + 1.

OpenStudy (yanasidlinskiy):

\(\Huge\bf \color{yellow}{Welcome~to~OpenStudy!!}\hspace{-310pt}\color{cyan}{Welcome~to~OpenStudy!!}\hspace{-307.1pt}\color{midnightblue}{Welcome~to~\color{purple}{Open}}\color{blue}{Study!!!}\)

OpenStudy (mathmale):

First things first. Your y = -2x^2 - 12x + 1 is a quadratic equation and has three constant coefficients: a, b and c. What are the values of a, b and c here? Next, what is the formula for the x-coordinate of the vertex of a parabola?

OpenStudy (anonymous):

first coordinate of the vertex of \[y=ax^2+bx+c\] is \(-\frac{b}{2a}\)

OpenStudy (mathmale):

rg: would you please answer my question(s): what are the values of a, b and c here?

OpenStudy (anonymous):

the values of a b and c are a=-2x^2 b=-12x and c=1

OpenStudy (mathmale):

Actually, rg, the coefficients are a=-2, b=-12 and c=1. I was asking you to find those coefficients, without the power-of-x terms. Now please take those values of a and b and substitute them into the formula for the x-coordinate of the vertex.

OpenStudy (mathmale):

See satellite73's post, above, for that formula.

OpenStudy (anonymous):

okay so it would be -12/2(-2) which is -12/-4 and that equals 3 right?

OpenStudy (mathmale):

\[x _{vertex}=\frac{ -b }{ 2a }\rightarrow x _{v}=\frac{ -(-12 ) }{ 2(-2) }\]=24/(-4) = ?

OpenStudy (anonymous):

thats -6

OpenStudy (mathmale):

But wouldn't -(-12) be +12? Wouldn't 2(-2) be -4?

OpenStudy (anonymous):

uhm yeah but then that would make it -3

OpenStudy (mathmale):

Yes, that's right. Now, if you want to find the vertex, take that \[x _{vertex}=-3\]and substitute it for x in your function, \[f(x)=-2x^2-12x+1\]to find the y-coordinate of the vertex. Last, write out the point that represents the vertex: It will look like (-3, ?? ).

OpenStudy (mathmale):

Sorry, but I'm getting off OpenStudy now. I believe you have enuf info with which to finish this problem. If not, you could always ask for help again. Good luck!